OpenStudy (anonymous):

im having a little trouble on this question 4y = 144 A: 16 B:-36 C:-16 D: 36

OpenStudy (alurahenderson):

Which ones do you think aren't the answer?

pooja195 (pooja195):

divide both sides by 4

jagr2713 (jagr2713):

Hey there @camacho01erick :D Ok what do you thin we do ?

pooja195 (pooja195):

This will isolate y :)

jagr2713 (jagr2713):

we are trying to get y by its self so we divide

jagr2713 (jagr2713):

So whats \[144\div 4\] @camacho01erick

OpenStudy (anonymous):

36

OpenStudy (alurahenderson):

Also remember positive x positive = positive and negative x positive = negative

jagr2713 (jagr2713):

So the answer is? @camacho01erick

OpenStudy (alurahenderson):

Good job!!!

OpenStudy (anonymous):

OOOOOH OK SO THE ANSWER IS 36

jagr2713 (jagr2713):

Correct \[ \huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color {green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid} {\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed} {\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color {magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color {ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\] if you want to know more about this site, Here is a tutorial about this site : http://openstudy.com/study#/updates/5525c0fce4b030ceb3f357e8

OpenStudy (anonymous):

thx guys

jagr2713 (jagr2713):

Anytime :D

OpenStudy (alurahenderson):

Sure:]

OpenStudy (anonymous):

thx guys you guys are the best :3

OpenStudy (alurahenderson):

If you need more help @ me :}}

OpenStudy (anonymous):

ok theres another one

OpenStudy (anonymous):

-1/8c=2

OpenStudy (alurahenderson):

Open a new question first:)

OpenStudy (anonymous):

oh ok